# Comment on “Galvano-rotational effect induced by electroweak interactions in pulsars”

### v1

@@ Abstract
In this comment, we obtain the corrected energy spectrum for Dvornikov’s paper, that is, a spectrum in which the matter term, given by $V_{L}$, is “tied” in both the radial and angular quantum numbers, given by $N$ and $J_{z}$. In particular, the error committed by Dvornikov arose from an incorrectly solved second-order differential equation. Besides, such an equation also had another error, i.e. since Dvornikov considered ultrarelativistic particles where $m\to 0$, it implies that all terms of this equation containing $m$ should vanish; however, that is not what happened. Therefore, solving in detail the corrected Dvornikov equation, we obtain the corrected spectrum, where we verified that for $J_{z}>0$ (positive angular momentum), $\sigma=+1$ (upper component of the spinor) and $V_{L}>0$, or $J_{z}<0$ (negative angular momentum), $\sigma=-1$ (lower component of the spinor) and $V_{L}<0$, we recover the particular spectrum of Dvornikov’s paper with $m^{2}=0$ (as it should be).
@@ I Introduction
In a paper published in the Journal of Cosmology and Astroparticle Physics (JCAP), entitled “Galvano-rotational effect induced by electroweak interactions in pulsars”, Dvornikov [2] studied the electroweakly interacting particles in rotating matter, where the existence of the electric current along the axis of the matter rotation was predicted in this system. According to him, this new Galvano-rotational effect is caused by the parity-violating interaction between massless charged particles in the rotating matter. To do such a study, he started with the exact solution of the Dirac equation for a fermion involved in the electroweak interaction in the rotating frame with constant angular velocity in cylindrical coordinates, where he used the vierbein vectors formalism to write the Dirac equation in a curved spacetime. Then, using the obtained solution (for $m\to 0$, i.e. ultrarelativistic neutrinos), he derived the induced electric current which turns out to flow along the rotation axis, and studied the possibility of the appearance of the Galvano-rotational effect in the dense matter of compact astrophysical objects. Later, he discussed the particular example of neutron and hypothetical quark stars, where it was shown that using this effect, one can expect the generation of toroidal magnetic fields comparable with poloidal ones in old millisecond pulsars. Besides, he also briefly discussed the generation of the magnetic helicity in these stars. Finally, he analyzed the possibility of applying the Galvano-rotational effect for the description of the asymmetric neutrino emission from a neutron star to explain pulsar kicks. In particular, this paper is well-written and addresses a very interesting topic involving Dirac neutrinos in rotating frames of reference and interacting with a background matter. Besides, the formalism used by him is also very important in the literature when working with Dirac fermions in curved space-times or rotating frames [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
However, analyzing in detail the energy spectrum in Dvornikov’s paper [2], the following question/doubt arose: Why does only a part of the Dvornikov energy spectrum depend on the angular quantum number $J_{z}$? (in which is given by $J_{z}\omega$, where $\omega$ is the angular velocity of the rotating frame). In other words, Why is the matter term in which the neutrino interacts, given by $V_{L}$, not also “tied” to the angular quantum number $J_{z}$? (in which is given by $4|V_{L}|N\omega$, where $N$ is the radial quantum number). To be more specific, the matter term in Dvornikov’s paper [2] that we refer to is that found inside a square root (which is a feature of any relativistic spectrum) and not that found outside, in which it is a free term (i.e. a term “tied” to nothing). In particular, such a question/doubt arose from the fact that Dvornikov [2] did not make it explicit/clear why $J_{z}$ also did not appear inside of the square root (or “tied” to $V_{L}$), such as occurs with $N$. In this way, this motivated us to look by the correct spectrum for the neutrino, that is, a spectrum in which the matter term is “tied” in both $N$ and $J_{z}$, and not only to $N$ as shown by Dvornikov [2]. Therefore, the goal of the present comment is to obtain the corrected energy spectrum for Dvornikov’s paper [2]. To achieve this, we use as a basis Refs. [3, 4, 7, 9, 10, 11, 12, 13, 14, 16, 17], where the Dirac equation was worked in curved space-times (with and without rotation). Besides, it is also important to mention that something similar to this has also been done by Dvornikov in another paper of his, but published in the JHEP [18] (i.e. the energy spectrum did not depend on $J_{z}$, whether inside or outside the square root). However, recently the corrected/complete energy spectrum of this paper was obtained by Ref. [19].
@@ II Quick review of the main steps that Dvornikov took to obtain the energy spectrum
$[i\gamma^{\mu}(x)\nabla_{\mu}-m]\psi=\gamma_{\mu}(x)\left\{\frac{V^{\mu}_{L}}{2}[1-\gamma^{5}(x)]+\frac{V^{\mu}_{R}}{2}[1+\gamma^{5}(x)]\right\}\psi,$ (1)
where $\gamma^{\mu}(x)=e^{\ \mu}_{a}\gamma^{a}$ and $\gamma_{\mu}(x)=e^{a}_{\ \mu}\gamma_{a}$ are the curved gamma matrices, $\gamma^{a}$ and $\gamma_{a}$ are the flat gamma matrices, $e^{\ \mu}_{a}$ is the vierbein and $e^{a}_{\ \mu}$ is its inverse (and satisfy the orthogonality condition $e^{a}_{\ \mu}e^{\ \mu}_{b}=\delta^{a}_{b}$), $\nabla_{\mu}=\partial_{\mu}+\Gamma_{\mu}$ is the covariant derivative, $\Gamma_{\mu}=-\frac{i}{4}\sigma^{ab}\omega_{ab\mu}=-\frac{i}{4}\sigma^{ab}e^{\ \nu}_{a}e_{b\nu;\mu}$ is the spin connection, $\sigma^{ab}=\frac{i}{2}[\gamma^{a},\gamma^{b}]$ are the generators of the Lorentz transformations in a locally Minkowskian frame, $\gamma^{5}(x)=-\frac{i}{4!}E^{\mu\nu\alpha\beta}\gamma_{\mu}(x)\gamma_{\nu}(x)\gamma_{\alpha}(x)\gamma_{\beta}(x)$ is the curved fifth gamma matrix, $E^{\mu\nu\alpha\beta}=\frac{1}{\sqrt{-g}}\varepsilon^{\mu\nu\alpha\beta}$ is the covariant antisymmetric tensor in curved space-time, $g$ is the determinant of the metric (i.e. $g=$ det$(g_{\mu\nu})$), $V^{\mu}_{L,R}$ are the effective potentials (with $V^{\mu}_{L}\neq V^{\mu}_{R}$), i.e. the last term of Eq. (1) is the background matter or simply the matter term, and $\psi$ is the four-component Dirac spinor, respectively.
@@
where $x^{\mu}=(t,r,\phi,z)$ is the position four-vector in cylindrical coordinates, $\omega$ is the constant angular velocity, and $g_{\mu\nu}$ is the metric tensor, respectively. In addition, one can check that the metric tensor in Eq. (2) can be diagonalized (via $\eta_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}g_{\mu\nu}$) if we use the following vierbein vectors [2]:
@@
where $\eta_{ab}$=diag(+1,1,1,1)$1$(+1,-1,-1,-1)( + 1 , - 1 , - 1 , - 1 ) is the metric in a locally Minkowskian frame. With this, the only non-zero component of the connection one-form $\omega_{ab}=\omega_{ab\mu}dx^{\mu}$ is given as follows [2] (originally obtained by Ref. [3]):
@@
From this, Dvornikov [2] rewrote Eq. (1) in the following equation (with $V^{\mu}_{L,R}=(V^{0}_{L,R},0,0,0)=(V_{L,R},0,0,0)$ and $\gamma^{5}(x)=\gamma^{5}=\gamma_{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$):
@@
[…]
@@
$+\left[\left(V_{V}\omega+\frac{J_{z}}{r^{2}}\right)\Sigma_{3}-p^{2}_{z}+V_{A}^{2}(1-\omega^{2}r^{2})\gamma^{5}+2mV_{A}(\gamma^{0}-\omega r\gamma^{2})\gamma^{5}-m^{2}\right]\Phi=0,$ (11)
@@
So, according to Dvornikov [2], the solution of Eq. (II) can be found for ultrarelativistic particles (i.e. in the limit $m\to 0$). With this, it is possible write $\Phi$ in the form $\Phi=\upsilon\varphi$, where $\upsilon$ is a constant spinor (and satisfies $\Sigma_{3}\upsilon=\sigma\upsilon$ and $\gamma^{5}\upsilon=\chi\upsilon$, where $\sigma=\chi=\pm 1$) and $\varphi=\varphi(r)$ is a scalar function [2]. Besides, Dvornikov [2] first studied left particles where $(1+\gamma^{5})\psi=0$, in which corresponds $\chi=+1$ (the case of right particles with $\chi=-1$ can be studied analogously). In this way, using a new variable in Eq. (II) given by $\rho=|V_{L}|\omega r^{2}$ (i.e. making a change of variable), Dvornikov [2] obtained the following equation for $\varphi_{\sigma}$:
@@
According to the spectrum above, we clearly see that the matter term (given by $V_{L}$) inside the square root is not “tied” to the angular quantum number $J_{z}$ (or $l$), but only to the radial quantum number $N$ (given by $4|V_{L}|N\omega$). That is, only a part of the Dvornikov energy spectrum depends on $J_{z}$, which is the second term of the spectrum and is “tied” to $\omega$ (given by $J_{z}\omega$). Besides, we also think very strange the presence of the (quadratic) mass term in the spectrum (given by $m^{2}$) since Dvornikov [2] said that the solution of Eq. (II) can be found for ultrarelativistic particles (i.e. $m\to 0$). In other words, this spectrum is incorrect (due to the presence of $m^{2}$ and the absence of $J_{z}$ “tied” to the matter term $V_{L}$ inside the square root).
III The corrected energy spectrum for Dvornikov’s paper
Before we obtain the corrected spectrum for Dvornikov’s paper, we first need to correct Eq. (12). So, as stated by Dvornikov [2], the solution of Eq. (II) can be found for ultrarelativistic particles ($m\to 0$), which implies that all terms of Eq. (12) containing $m$ should vanish (or be zero); however, that is not what happened since there is the presence of $m^{2}$ in the equation. In other words, Dvornikov [2] zeroed one term, given by $2mV_{A}(\gamma^{0}-\omega r\gamma^{2})\gamma^{5}$, and kept another, given by $m^{2}$ (which could not have happened, i.e. it is wrong). Besides, is also important to mention here that the condition for we have only left particles in the Dirac equation (1) is if $V^{\mu}_{R}=0$ (or $V_{R}=0$) [18], and do not for $(1+\gamma^{5})\psi=0$, as stated by Dvornikov [2]. Consequently, this implies that: $V_{V,A}=V_{L}/2\neq 0$ (Dvornikov used this to arrive at (12)). In short, Eq. (12) is only really correct if the quadratic mass term is also zero ($m^{2}=0$); consequently, the equation must not have any terms containing $m$ (in particular, that should be the case since Dvornikov did an earlier paper, given by [18], in which the equation does not contain $m$).
So, making $m^{2}=0$ as well as excluding the terms ±$\pm\frac{1}{2}\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right)$ in Eq. (12), that is, there is no need to use such terms, we have the following equation (with $\partial_{\rho}\to\frac{d}{d\rho}$):
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\frac{d}{d\rho}-\frac{\gamma^{2}}{4\rho}-\frac{\rho}{4}+\beta\right]\varphi_{\sigma}(\rho)=0,$ (15)
where we define:

### v2

@@ Abstract
In this comment, we show that the fundamental equation worked by Dvornikov in his paper is incorrect, which is given by the Dirac equation for a neutrino interacting with matter in a rotating frame (or curved space-time). In particular, Dvornikov incorrectly wrote/defined the interaction term (effective potentials or background matter) in a curved space-time. Consequently, one of his main results (if not the main one), namely the relativistic energy spectrum, is also incorrect (actually partially incorrect). So, starting from the correct Dirac equation with effective potentials in a curved space-time, we obtain the correct energy spectrum for a neutrino interacting with matter in a rotating frame. We observe that the square root of this spectrum is very different from that obtained by Dvornikov. In fact, in our case, the spectrum was obtained from a Bessel equation subject to a hard-wall confining potential, while the Dvornikov spectrum was obtained from an associated Laguerre equation. However, we note that the only correct thing in the Dvornikov spectrum are the terms outside the square root, which agree with those of the correct spectrum.
@@ I Introduction
In a paper published in the Journal of Cosmology and Astroparticle Physics (JCAP), entitled “Galvano-rotational effect induced by electroweak interactions in pulsars”, Dvornikov [2] studied the electroweakly interacting particles in rotating matter, where the existence of the electric current along the axis of the matter rotation was predicted in this system. According to Dvornikov [2], this new Galvano-rotational effect (GRE) is caused by the parity-violating interaction between massless charged particles in the rotating matter. To do such a study, Dvornikov [2] worked with the fundamental equation of relativistic quantum mechanics for spin-1/2 particles in curved space-times, that is, with the Dirac equation for a neutrino interacting with matter in a rotating frame (with constant angular velocity and in cylindrical coordinates), where the formalism used to write this equation in a curved space-time (or rotating frame) was the well-known vierbein vectors formalism (of general relativity of Einstein). Therefore, this equation includes the noninertial effects generated by a uniformly rotating frame. So, by defining a new spinor for the system, Dvornikov [2] transformed the Dirac equation into a second-order (ordinary) differential equation. Considering the limit $m\to 0$ (i.e. ultrarelativistic or massless neutrinos), Dvornikov [2] solved the differential equation and obtained the relativistic energy spectrum (or quantized energy eigenvalues) for the neutrino, where such a result was one of the most important (if not the greatest) of his paper. In fact, it is through this spectrum (with some conditions) that Dvornikov [2] obtained the induced electric current (hydrodynamic current) along the rotation axis, where this phenomenon was called by him the new galvano-rotational effect. In particular, this paper is well-written and addresses a very interesting topic involving Dirac neutrinos interacting with matter in a rotating frame. As for the formalism used, it is also very important in the literature when working with Dirac fermions in curved space-times [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
However, analyzing in detail a paper recently published in Physical Review D (PRD) [20], we verified that the Dirac equation worked/used by Dvornikov [2] in his paper is incorrect, that is, the Dirac equation for a neutrino interacting with matter in a rotating frame (or curved space-time) was written incorrectly. To be even more specific, Dvornikov [2] incorrectly wrote/defined the interaction term (effective potentials or background matter) in a curved space-time. Consequently, one of his main results (if not the main one), namely the relativistic energy spectrum (or quantized energy eigenvalues), is also incorrect. Therefore, using Ref. [20] as a basis (as well as other related papers), the present comment has as its goal to show in detail the correct Dirac equation with effective potentials in a curved space-time and, consequently, to obtain the correct energy spectrum for a neutrino interacting with matter in a rotating frame. To achieve the quantization of energy (or relativistic bound states), we will consider that our second-order differential equation is subject to a hard-wall confinement potential (analogous to the well-known finite/infinite square potential well). Without this potential, the quantization of the energy (or bound states) is not achieved. That is, unlike the Dvornikov differential equation [2], here we will obtain a Bessel equation.
@@ II Quick review of the main steps that Dvornikov took to obtain the energy spectrum
$i\gamma^{\mu}(x)\nabla_{\mu}-m]\psi=\gamma_{\mu}(x)\left\{\frac{V^{\mu}_{L}}{2}[1-\gamma^{5}(x)]+\frac{V^{\mu}_{R}}{2}[1+\gamma^{5}(x)]\right\}\psi,$ (1)
[…]
@@
where $x^{\mu}=(t,r,\phi,z)$ is the position four-vector in cylindrical coordinates, $\omega$ is the constant angular velocity, and $g_{\mu\nu}$ is the metric tensor, respectively. In addition, one can check that the metric tensor in Eq. (2) can be diagonalized (via $\eta_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}g_{\mu\nu}$) if we use the following vierbein vectors (or vierbeins) [2]:
@@
where $\eta_{ab}$=diag(+1,1,1,1)(+1,-1,-1,-1) is the metric in a locally Minkowskian frame. With this, the only non-zero component of the connection one-form $\omega_{ab}=\omega_{ab\mu}dx^{\mu}$ is given as follows [2] (originally obtained by Ref. [3]):
@@
From this, Dvornikov [2] rewrote Eq. (1) in the following equation (with $V^{\mu}_{L,R}=(V^{0}_{L,R},0,0,0)=(V_{L,R},0,0,0)$ and $\gamma^{5}(x)=\gamma^{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$, i.e. $V^{\mu}_{L,R}$ and $\gamma^{5}(x)$ do not depends on vierbein or metric [21]):
@@
$\gamma^{0}=\left(\begin{array}[]{cc}0&-1\\ -1&0\\ \end{array}\right),\ \ \gamma^{k}=\left(\begin{array}[]{cc}0&\sigma^{k}\\ -\sigma^{k}&\ 0\\ \end{array}\right),\ \ \gamma^{5}=\left(\begin{array}[]{cc}1&\ 0\\ 0&-1\\ \end{array}\right).$ (7)
@@
$+\left[\left(V_{V}\omega+\frac{J_{z}}{r^{2}}\right)\Sigma_{3}-p^{2}_{z}+V_{A}^{2}(1-\omega^{2}r^{2})+2mV_{A}(\gamma^{0}-\omega r\gamma^{2})\gamma^{5}-m^{2}\right]\Phi=0,$ (11)
@@
So, according to Dvornikov [2], the solution of Eq. (II) can be found for ultrarelativistic particles (i.e. in the limit $m\to 0$). With this, it is possible write $\Phi$ in the form $\Phi=\upsilon\varphi$, where $\upsilon$ is a constant spinor (and satisfies $\Sigma_{3}\upsilon=\sigma\upsilon$ and $\gamma^{5}\upsilon=\chi\upsilon$, where $\sigma=\chi=\pm 1$) and $\varphi=\varphi(r)$ is a scalar function [2]. Besides, Dvornikov [2] first studied left-handed particles where $(1+\gamma^{5})\psi=0$ or $V_{R}=0$ (with $V_{V,A}=V_{L}/2$), which corresponds to $\chi=+1$ (the case of right particles with $\chi=-1$ can be studied analogously). In this way, using a new variable in Eq. (II) given by $\rho=|V_{L}|\omega r^{2}$ (i.e. making a change of variable), Dvornikov [2] obtained the following equation for $\varphi_{\sigma}$:
@@
[…]
III The correct Dirac equation with effective potentials in a curved space-time for a neutrino interacting with matter in a rotating frame
Here, let us get the correct Dirac equation with effective potentials in a curved space-time for a neutrino interacting with matter in a rotating frame. So, according to Ref. [20], the Dirac action for a neutrino interacting with matter in a curved space-time is given by
$S_{int}=\int d^{4}x\sqrt{-g}\mathcal{L}_{int},$ (15)
where $\mathcal{L}_{int}$ is the interaction Lagrangian density (due to the charged current interaction effects), given as follows
$\mathcal{L}_{int}=-\frac{G_{F}}{\sqrt{2}}J_{\mu}\bar{\nu}_{e}\gamma^{\mu}(1-\gamma^{5})\nu_{e},\ \ (\mu=t,r,\tilde{\theta},\varphi),$ (16)
being $G_{F}$ the Fermi coupling constant, $J_{\mu}=\bar{\psi}_{e}\gamma_{\mu}(1-\gamma^{5})\psi_{e}$ is the charged current due to the electron field $\psi_{e}$ (i.e. Dirac spinor of the electron), $\gamma_{\mu}=\gamma_{\mu}(x)=e_{\ \mu}^{b}\gamma_{b}$ or $\gamma^{\mu}=\gamma^{\mu}(x)=e^{\mu}_{\ a}\gamma^{a}$ are curved gamma matrices, and $\nu_{e}$ is the electron neutrino field (i.e. Dirac spinor of the electron neutrino). Indeed, using the standard prescription (“standard recipe’) widely used in the literature to transform the flat Dirac equation into the curved Dirac equation, i.g. $d^{4}x\to d^{4}x\sqrt{-g}$, $\gamma^{\mu}\to\gamma^{\mu}=e^{\mu}_{\ a}\gamma^{a}$ and $\partial_{\mu}\to\nabla_{\mu}=\partial_{\mu}+\Gamma_{\mu}$ [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 21, 32], into flat interaction Lagrangian given in Refs. [33, 34, 35, 36] (with $\mathcal{H}_{int}=-\mathcal{L}_{int}$), we get exactly (16) (or (15)). Therefore, the curved interaction Lagrangian (16) (or curved action (15)) is indeed correct.
So, according to Ref. [20], in a static and spherically symmetric space-time (but can be cylindrical too), we can approximate the background electron current as $J_{\mu}=n_{e}e^{0}_{\ \mu}$, where $n_{e}=n_{e}(x)$ is the electron number density or simply the electron density. Consequently, the interaction Lagrangian (16) reduces to [20] (we believe this reference forgot the minus sign)
$\mathcal{L}_{int}=-\frac{G_{F}}{\sqrt{2}}J_{t}\bar{\nu}_{e}\gamma^{t}(1-\gamma^{5})\nu_{e}=-\frac{G_{F}}{\sqrt{2}}n_{e}e^{0}_{\ t}\bar{\nu}_{e}e^{t}_{\ 0}\gamma^{0}(1-\gamma^{5})\nu_{e}=-\frac{G_{F}}{\sqrt{2}}n_{e}\bar{\nu}_{e}\gamma^{0}(1-\gamma^{5})\nu_{e}=-(\sqrt{2}G_{F}n_{e})[\bar{\nu}_{e}\gamma^{0}P_{L}\nu_{e}],$ (17)
where $P_{L}=\frac{1}{2}(1-\gamma^{5})$ is the left-handed projection operator, and we use the fact that $e^{0}_{\ t}e^{t}_{\ 0}=\delta^{t}_{t}=\delta^{0}_{0}=1$. Therefore, we see that $\mathcal{L}_{int}$ does not contain any explicit dependence on vierbein or space-time metric (i.e. $\mathcal{L}_{int}^{curved}=\mathcal{L}_{int}^{flat}$) [20] (however, the metric dependence is contained in the term $\sqrt{-g}$ in (15) [20]).
Now, according to Ref. [20], the dynamics of a free fermion in a curved space-time is governed by the following action
$S_{\psi}=\int d^{4}x\sqrt{-g}\mathcal{L}_{\psi},$ (18)
where $\mathcal{L}_{\psi}$ is the free Dirac Lagrangian density (or free Lagrangian density), given as follows
$\mathcal{L}_{\psi}=-\bar{\psi}[i\gamma^{\mu}D_{\mu}+m]\psi,$ (19)
being $D_{\mu}=\nabla_{\mu}=\partial_{\mu}+\Gamma_{\mu}$ the (spin) covariant derivative, $\Gamma_{\mu}$ is the spin connection, and $m$ is the mass term (of the fermion). However, according to Ref. [32] (page 212), the minus sign in (19) is in the wrong place, i.e. the minus sign should be in the mass term, such as $\mathcal{L}_{\psi}=\bar{\psi}[i\gamma^{\mu}D_{\mu}-m]\psi$ (we believe this is the correct Lagrangian). So, doing $\bar{\nu}_{e}\to\bar{\psi}$ and $\nu_{e}\to\psi$ (i.e. consider only a single neutrino flavor where the mixing angle is null/zero), implies that the total/complete Lagrangian for the neutrino (electron neutrino, actually) will be given by (with the negative signs already corrected)
$\mathcal{L}_{total}=\bar{\psi}\left[i\gamma^{\mu}D_{\mu}-m-(\sqrt{2}G_{F}n_{e})\gamma^{0}P_{L}\right]\psi.$ (20)
Therefore, using the following Euler-Lagrange equation [37]
$\partial_{\mu}\frac{\partial\mathcal{L}_{total}}{\partial(\partial_{\mu}\bar{\psi})}-\frac{\partial\mathcal{L}_{total}}{\partial_{\mu}\bar{\psi}}=0,$ (21)
we can obtain the following Dirac equation for a (left-handed) neutrino interacting with matter in a curved space-time
$\left[i\gamma^{\mu}D_{\mu}-m-(\sqrt{2}G_{F}n_{e})\gamma^{0}P_{L}\right]\psi=0,$ (22)
or yet
$\left[i\gamma^{\mu}(x)D_{\mu}-m\right]\psi=\gamma_{\mu}(x)V^{\mu}_{L}(x)P_{L}\psi,$ (23)
where we define $V^{\mu}_{L}(x)=e^{\ \mu}_{a}V^{a}_{L}$ (left-handed effective potential), with $V^{L}_{a}=(V^{0}_{L},0,0,0)=(\sqrt{2}G_{F}n_{e},0,0,0)$, and $\gamma_{\mu}=e^{a}_{\ \mu}\gamma_{a}$ (in particular, a more general case where $n_{e}\to n_{e}-n_{n}/2$, being $n_{n}$ the neutrons density, is give by [35, 36]). However, to also include a right-handed neutrino in the equation, simply do: $V^{\mu}_{L}(x)P_{L}(x)\to V^{\mu}_{L}P_{L}(x)+V^{\mu}_{R}(x)P_{R}$, where $V^{\mu}_{R}(x)=e^{\ \mu}_{a}V^{a}_{R}$ (with $V^{a}_{R}=(V^{0}_{R},0,0,0)$) would be the right-handed effective potential (with $V^{\mu}_{R}(x)\neq V^{\mu}_{L}(x)$ or $V^{0}_{R}\neq V^{0}_{L}$), being $P_{R}=\frac{1}{2}(1+\gamma^{5})$ the right-handed protection operator (in particular, considering that $n_{e}$ and $n_{n}$ do not depend on the coordinates, i.e. are uniform quantities, implies that theses effective potentials are the same worked by Dvornikov [2] in his Table 1). Therefore, the correct Dirac equation for Dvornikov’s paper [2] (Dirac equation with effective potentials in a curved space-time) should be given by
$i\gamma^{\mu}(x)\nabla_{\mu}(x)-m]\psi=\gamma_{\mu}(x)\left\{\frac{V^{\mu}_{L}(x)}{2}[1-\gamma^{5}(x)]+\frac{V^{\mu}_{R}(x)}{2}[1+\gamma^{5}(x)]\right\}\psi,$ (24)
where both $\gamma_{\mu}(x)V^{\mu}_{L}(x)$ and $\gamma_{\mu}(x)V^{\mu}_{R}(x)$ (and $\gamma^{5}(x)$) do not depend on vierbein or metric ($\gamma_{\mu}(x)V^{\mu}_{L/R}(x)=\gamma_{0}V^{0}_{L/R}$). Before proceeding, it is important to mention that although Dvornikov also worked with Eq. (1) (incorrect equation) in another paper (given by Ref. [18]), the result still remained correct (but incomplete, as shown by Ref. [19]) since the slow rotation regime was considered (which implies that $e_{\ t}^{0}\approx 1$). In this case (slow rotation), it makes no difference whether $g^{\mu}$ depends on $e^{\mu}_{a}$ or not (again, the correct thing is to depend on $e^{\mu}_{a}$). However, we will see that here it makes a significant difference, that is, unlike [18], here, Dvornikov [2] chose another way of including rotation in the system. Consequently, we will see that the Dirac equation (24) will generate another type of second-order differential equation (which is different from Refs. [18, 2], where an associated Laguerre equation was generated).
So, as a consequence of Eq. (24), Eq. (5) should be written as
$\mathcal{D}\psi=\gamma^{0}(V_{V}-V_{A}\gamma^{5})\psi,$ (25)
whereas (9) should be
$\left[\gamma^{a}\tilde{Q}_{a}-m+\tilde{V}\right]\psi_{r}=0,$ (26)